Solve the system by substitution:
4x - 12y = 5
-x + 3y = -1
Thanks for any help you can give me.

Answer by ptaylor(2198)

Solve the system by substitution
(1) 4x - 12y = 5
(2) -x + 3y = -1
Thanks for any help you can give me.

(1) 4x - 12y = 5
(2) -x + 3y = -1
subtract 3 from both sides of (1) and multiply both sides by -1
(1) 4x - 12y = 5
(2) x=1+3y
Substitute x=1+3y into (1)
4(1+3y)-12y=5 clear parens
4+12y-12y=5----------------------------------BLOWS APART
Lets look at the slope of Eq (1) and (2)
y=mx+b, where m is the slope
(1) y=(4/12)x-5/12
(2) y=(1/3)x-1

These lines have the same slope. That means that they are parallel.
There's no solution-------------------answer

Hope this helps---------------ptaylor